A 20.0 cm long section of wire is placed on top of a sensitive electronic scale. A second wire is suspended directly above and parallel to the first wire. The distance between the wires is2.00 cm. The wires are part of a continuous circuit with a current (I) flowing in both wires in opposite directions. When the current is off, the scale reads the mass of the segment of wire, when the

current is on, the scale reading increases by 5.00 mg. How much current is running through the wires? Consider the top wire to be infinitely long.
Answer #1

when current is off

W = M*g


when current is on W' = F + Mg

W' = (M + 0.005 )*g = Mg + 0.005*g


therfore

F = 0.005*g


but F = magnetic force experienced by wire = uo*I*I*L/(2*pi*r)


uo*I*I*L/(2*pi*r) = 0.005*g


4*pi*10^-7*I^2*0.2/(2*pi*0.02) = 0.005*9.81


I = 156.6 A